2. Using calculus, find the mean and variance of an exponential distribution with a probability density function of f(x)-Aet, L? (Give a proof What is A in terms of
Answer #1

for above to be vald: \int_{0}^{\infty } f(x) dx =1

\int_{0}^{\infty } f(x) dx=\int_{0}^{\infty } Ae-x/L dx =-ALe-x/L |\infty0 =-AL*(0-1) =AL=1

A=1/L

hence mean =E(X)=\int_{0}^{\infty } xf(x) dx =\int_{0}^{\infty } (x/L)e-x/L dx =-(xe-x/L-Le-x/L) |\infty0 =-(0-L) =L

E(X2)=\int_{0}^{\infty } (x2/L)e-x/L dx =-(x2e-x/L-2xLe-x/L-2L2e-x/L) |\infty0 =-(0-2L2) =L2

therefore Variance =E(X2)-(E(X))2 =2L2-L2 =L2